# Supervised Score-Based Modeling by Gradient Boosting

### v1

@@ Introduction
CIFAR Accuracy 90.93 $\pm$ 0.02 90.91 $\pm$ 0.00 90.99 $\pm$ 0.00
-10 Time 50.30 $\pm$ 0.30 19.63 $\pm$ 1.21 15.43 $\pm$ 1.12
CIFAR Accuracy 71.42 $\pm$ 0.01 71.51 $\pm$ 0.00 71.38 $\pm$ 0.00
-100 Time 90.58 $\pm$ 0.12 32.30 $\pm$ 0.60 27.16 $\pm$ 0.19
@@
Table 4: Accuracy (in %) and inference time for one batch (in second (s)) by CARD and SSM on CIFAR-10 and CIFAR-100.
Experiment
Toy examples
To demonstrate the effectiveness of SSM and all improved Techniques, we first perform experiments on 5 selected toy examples (linear regression, quadratic regression, log-log linear regression, log-log cubic regression, and sinusoidal regression) proposed in (Han, Zheng, and Zhou 2022). We add unbiased normal distribution noise to the original models to obtain noisy data for training, increasing model complexity. We aim to predict accurate regression results for pure input data in the inference process. To evaluate the impact of various technologies, we have designed an ablation experiment that considers three variables: the use of original Langevin dynamics, the implementation of the fast sampling technique as described in Technique 2, and the application of different coefficients $\lambda(\sigma_{i})$ in the loss function, where L1 and L2 represent the degree of noise levels of 1 and 2, respectively. The experiment results with the Root Mean Squared Error (RMSE) and inference time about different model settings are shown in Table 1 and 2, respectively. The scatter plots for 5 examples of both true and generated data following the first setting, i.e., L1, w/o noise fast, are shown in Figure 1. We observe that while perturbed data is used, SSM can still fit the regression equation of the original data. The proposed fast sampling Technique 2 significantly reduces the inference time for all tasks. Additionally, since the fast sampling technique ensures the error bound of the current estimated state, it can even achieve the best performance on tasks except for log-log linear regression. Using the noise-free Langevin equation, i.e., w/o noise, performs better on most tasks, indicating that the selection of different coefficients also affects the performance of the model.
Regression
We further evaluate our model on 10 UCI regression tasks (Dua and Graff 2017). We employ the same experimental settings as those used in the CARD model (Han, Zheng, and Zhou 2022). We compare our model with NGboost, CARD, DBT, and several Bayesian neural network frameworks considered in (Han, Zheng, and Zhou 2022), including MC Dropout (Gal and Ghahramani 2016), Deep Ensembles (Lakshminarayanan, Pritzel, and Blundell 2017), and GCDS (Zhou et al. 2023). The experiment results with RMSE are shown in Table 3. It is noticed that SSM outperforms existing approaches in RMSE. SSM (L2) obtains the best results in 8 out of 10 datasets. Although the RMSE of SSM (L1) is larger, it has a smaller standard deviation, which means the network is more stable. In addition, probabilistic models such as CARD predict the probability of distribution, while SSM predicts the points directly. This also leads to a lower standard deviation of SSM on most tasks.
Classification
Conclusion
In this paper, we proposed Supervised Score-based Model (SSM), score-based gradient boosting model using denoising score matching for supervised learning. First, we analyzed the connection between SSM and GBM and showed that the iterative equation of GBM can be converted to a noise-free Langevin equation. Then, we provided a theoretical analysis on learning and sampling for SSM to balance inference time and prediction accuracy with noise-free Langevin sampling. Furthermore, we showed the effectiveness of the proposed techniques on selected toy examples. Lastly, we compared SSM with existing models, including NGboost, CARD, DBT, and Bayesian neural network-based models, for several regression and classification tasks. The experimental results show that our model outperforms in both accuracy and inference time.

### v2

@@ Appendix B Experimental details
All 90.95% 90.91% 90.99%
1 91.00% 91.00% 90.70%
2 96.00% 96.10% 96.10%
3 87.30% 88.20% 88.30%
4 81.90% 81.30% 81.40%
5 93.30% 92.40% 94.10%
6 84.70% 85.50% 84.80%
7 94.20% 94.50% 94.50%
8 92.80% 92.40% 92.30%
9 95.30% 94.90% 95.00%
10 93.00% 92.80% 92.70%
# best 5 4 4
Table 14: Comparison of the accuracy of SSM and CARD in each category of CIFAR-10 classification task.
Model Acc. by $f_{\phi}$ CARD SSM (L1) SSM (L2)
Accuracy 71.37% 71.42 $\pm$ 0.01% 71.51 $\pm$ 0.00% 71.38 $\pm$ 0.00%
Time - 90.58 $\pm$ 0.12 32.30 $\pm$ 0.60 27.16 $\pm$ 0.19
@@
Table 15: Accuracy and mean inference time for one batch (in second (s)) by CARD and SSM on CIFAR-100. $f_{\phi}$ is the pre-trained base classifier.
Comparison with GBM
In this subsection, we elaborate on the similarities and differences between SSM and GBM methods, including NGboost and DBT.
From an algorithmic perspective, obtaining the score function, i.e., log-likelihood estimation, is generally challenging, making it difficult to obtain the maximum likelihood estimate in traditional GBM, including NGboost. Accordingly, to solve this issue, our SSM approach utilizes denoising score matching for score estimation, which has been proven on a variety of tasks that better estimation results can be achieved.
From the perspective of network structure, we train a unified model for the denoising process, like most score-based generative models. Compared to using multiple weak learners in GBM, utilizing a single model for gradient boosting can achieve parameter sharing, helping the network learn more general and powerful feature representations throughout the denoising process, and improve training efficiency.
Moreover, compared to DBT, which is also a score-based gradient boosting model, SSM utilizes the inference process in the same way as the original score-based generative model instead of a decision tree structure. Although using a decision tree model can enhance the interpretability of the model, the output process of the model is more dependent on the number of leaves. When the number of leaves is large, i.e., the denoising timesteps are long, the DBT model based on the decision tree cannot be effectively accelerated using the acceleration algorithm, including DDIM, resulting in a relatively long inference time for the model, which is also one of the advantages of SSM.
Additionally, in the traditional GBM model, the weak learners only use the original input, which leads to an unclear relationship between different learners. By leveraging denoising score matching, the proposed SSM has two advantages:
  • The previous denoising result is always used as input in the next denoising steps, which ensures the Markov property and enhances the correlation in denoising process.
  • Different noise levels ensure the focus of model learning and improve learning efficiency. Moreover, the performance of GBM model is limited by the settings of the hyperparameter.
In this paper, we discussed the setting of hyperparameters in detail and combined ablation experiments to ensure the effectiveness of our method in Section 4.
Comparison with CARD
Lastly, we make a detailed comparison between SSM and CARD and provide an insight analysis on their differences.
First, SSM is a score-based gradient boosting model, and CARD is a diffusion-based model. Although the two models are somewhat related, differences in design motivations lead to differences in the training and inference processes. The score-based model aims to estimate the gradient of unknown data distribution via score matching. The added noises are to facilitate estimation, where more noises can ensure that the gradients of more points in the space can be trained. Additionally, via the Langevin dynamics, we can approximate an unknown distribution and generate data from that distribution (Song and Ermon 2019). On the other hand, the diffusion-based model makes the unknown distribution converge to the standard normal distribution by adding noise. By learning the added noise, we can denoise the noise to the original distribution step by step, generating new data (Ho, Jain, and Abbeel 2020).
This difference leads to the difference in the number denoising steps between the two methods. The score-based model needs an enough number of steps to make the Langevin dynamics converge. Since it estimates the noise added at each step, the diffusion-based model can only denoise according to the specified number of steps. Therefore, the proposed model SSM can utilize score function as the current estimated error distance to ensure the convergence of the Langevin dynamics, thus reducing the number of denoising steps, while CARD cannot achieve the same result.
Figure 2: One batch prediction results corresponding to different denoising steps on UCI-Yacht task by SSM (top) and CARD (bottom). $n$ represents the number of denoising steps. The red and blue points represent the true values and the prediction results generated by 1000 samples respectively.
Another important difference is the different goals of model prediction. Since SSM removes the noise term in the Langevin dynamics, each step of denoising moves towards the maximum likelihood estimation point, i.e., the prediction of the input data. Even if we start denoising from different initial points, the convergence will always be concentrated in a small range of the target. Relatively, CARD and other Bayesian neural network frameworks are probabilistic prediction models, whose aim is to the distribution that the outcome may follow. To illustrate this clearly, the prediction results corresponding to different denoising steps on UCI-Yacht task by SSM and CARD are shown in Figure 2. It is noticed that for 1000 prediction points generated by the same input, a more compact prediction result can be obtained by SSM, which also shows that SSM is a more deterministic model. Moreover, the SSM figure is plotted at the step when SSM switches to the next noise level for denoising. It is noticed that under the guarantee of error estimation, SSM can indeed ensure that the current state converges within the error range of the target point.
In summary, SSM is a relatively deterministic prediction model. Additionally, since the inference process is constantly shortening the distance between the current state and the target state, SSM can ensure the convergence of the model based on the error and shorten the inference time.
Limitations and Future Work
In this section, we discuss the limitations of SSM and our future work.
First, we apply Technique 1 by removing the noise term in the Langevin dynamics to make the model deterministic. Even though SSM outperforms the probabilistic models including CARD on several UCI tasks and CIFAR datasets, its deterministic prediction ability makes it possible to perform poorly when dealing with some tasks where the test set is too different from the training set, such as out of distribution dataset (Yang et al. 2021). For this type of task, probabilistic models can handle it more effectively because they predict probability distributions rather than deterministic values (Ren et al. 2019). Second, according to the average number of denoising steps shown in Table 11, we notice that the upper step number bound approximated by Technique 2 is usually smaller than the actual value. This is because we do not consider the network error when computing the end-signals, which may lead to a large gap between the estimated number of steps and the actual number, making it impossible to guarantee the completion time in practical applications. Additionally, due to the network error, the condition in Alg. 2 line 6 cannot always be satisfied, which will cause the current number of denoising steps to reach the given limited number. As a reuslt, there is no guarantee of error bounds in the subsequent denoising processes, leading to a decrease in prediction accuracy. Finally, We use the average loss of the network on the training set as the error of the score network used in the Technique 3, which may also lead to an inaccurate steps number approximation. Besides, when the lower bound computed by Technique 3 is small, this affects Techniques 4 whose aim is to train more on the last noise level.
In our future work, we plan to incorporate uncertain items into the SSM, such as in the last denoising stage, to enhance its ability to predict the probability distribution of the target value instead of the target value itself. Moreover, we will consider more accurate estimates from the previous period to minimize the impact of network errors on the estimation results. Finally, Our estimates of inference time and error are from a numerical perspective rather than a distributional perspective, which are not applicable to the original score-based generative model. Therefore, we will consider modifying our technique to accommodate score-based generative models.